JAIDENGYMNASTIC:

Find the quotient of 1 1/5 and 3 4/5

JAIDENGYMNASTIC:

wait

JAIDENGYMNASTIC:

im edeting it

Acolgrove:

kay

JAIDENGYMNASTIC:

ok i got it

JAIDENGYMNASTIC:

so now i will screen shot the options

JAIDENGYMNASTIC:

1 attachment
JAIDENGYMNASTIC:

there ya go

Acolgrove:

do you try the math and if u did what did u get

Acolgrove:

and i got a different awnser than that

mxddi3:

i got the same answer that you chose in the picture

JAIDENGYMNASTIC:

no i did not mean to click on it i have to click on somthing for me to answer the question

JAIDENGYMNASTIC:

it is a random choice

mxddi3:

oh lol, well what did YOU do first to solve this? or do you know what to do first?

JAIDENGYMNASTIC:

i dont i forgot how to divide mixed numbers

JAIDENGYMNASTIC:

well dont you have to turn it into a fraction first?

mxddi3:

ok, so first turn your mixed numbers into improper fractions, yes

JAIDENGYMNASTIC:

ok so 6/5

JAIDENGYMNASTIC:

?

mxddi3:

and? whats the other one

JAIDENGYMNASTIC:

uhhh let me see

JAIDENGYMNASTIC:

18/5

mxddi3:

close, 3x5 is 15 and then 15 + 4 is 19, so 19/5

JAIDENGYMNASTIC:

OHHH NO well dang

mxddi3:

then, do u know the next step?

JAIDENGYMNASTIC:

uh well u need to divide so find a way to .... OH WAIT can we do kfc with this

JAIDENGYMNASTIC:

?

mxddi3:

mhmm, yes. so we will then have \[\frac{ 6 }{ 5 } *\frac{ 5 }{ 19 }\]

mxddi3:

do you know how to cross multiply?

JAIDENGYMNASTIC:

yes

mxddi3:

i meant cross cancel, my bad. but the 5's can cancel out and give us \[\frac{ 6 }{ 1 } * \frac{ 1 }{ 19 }\] what does that give us?

JAIDENGYMNASTIC:

30/95

mxddi3:

ahh, that can be simplified. you can do it that way or you can just do it as i set it up 6x1= ? 1x19=?

JAIDENGYMNASTIC:

6/19

mxddi3:

yes, that's it. good job!

JAIDENGYMNASTIC:

thanks you

mxddi3:

of course!